# Threats, Attacks, and Defenses in Machine Unlearning: A Survey

### v1

@@ V-C Proof of Unlearning
As mentioned previously, membership inference attacks serve to identify information leakage within unlearning systems, while backdoor attacks aim to implant backdoors in models, leading to the models incorrectly classifying certain inputs when specific triggers are present. Viewing from a different angle, in a machine unlearning setting, if unlearned data show no membership in the model post-unlearning, or if backdoors within the unlearned data are undetectable, it is logical to conclude that the unlearned data has been effectively removed from the model with high probability. Based on this rationale, many studies on unlearning methodologies have employed the standard form or variants of inference attacks [93, 94, 95, 96, 97, 98, 99] and backdoor attacks [100, 28, 101, 102, 103, 104] as tools serving as proof of unlearning. Some other studies utilize adversarial attacks to detect the insufficiency of unlearning procedures [105, 106, 70, 71]. This list represents only a selection of studies, as the use of attacks to demonstrate unlearning has been thoroughly reviewed in prior literature. For those seeking more in-depth information on this field, we recommend referring to [17, 10, 15] for comprehensive details.
@@ VII Conclusions
This survey delves into machine unlearning, focusing on threats and attacks against these systems and the defensive strategies to mitigate them. It provides a detailed analysis of methodologies and creates a comprehensive taxonomy based on their threat models. Furthermore, the survey investigates how unlearning can act as a defense in different scenarios and how attacks can serve as effective tools for testing and improving unlearning systems. Lastly, this work identifies challenges and outlines future directions for improving the safety, reliability, and privacy compliance of machine unlearning, emphasizing the urgent necessity for future research.

### v2

@@ V-C Proof of Unlearning
As mentioned previously, membership inference attacks serve to identify information leakage within unlearning systems, while backdoor attacks aim to implant backdoors in models, leading to the models incorrectly classifying certain inputs when specific triggers are present. Viewing from a different angle, in a machine unlearning setting, if unlearned data show no membership in the model post-unlearning, or if backdoors within the unlearned data are undetectable, it is logical to conclude that the unlearned data has been effectively removed from the model with high probability. Based on this rationale, many studies on unlearning methodologies have employed the standard form or variants of inference attacks [100, 101, 102, 103, 104, 105, 106] and backdoor attacks [107, 31, 108, 109, 110, 111] as tools serving as proof of unlearning. Some other studies utilize adversarial attacks to detect the insufficiency of unlearning procedures [112, 113, 78, 79]. This list represents only a selection of studies, as the use of attacks to demonstrate unlearning has been thoroughly reviewed in prior literature. For those seeking more in-depth information on this field, we recommend referring to [18, 10, 16, 46, 47] for comprehensive details.
@@ VII Conclusions
This survey examines machine unlearning as a vital component of efforts towards achieving safe AI, concentrating on threats and attacks against these systems, alongside the defensive strategies employed to counteract them. It provides a detailed analysis of methodologies and creates a comprehensive taxonomy based on their threat models. Furthermore, the survey investigates how unlearning can act as a defense in different scenarios and how attacks can serve as effective tools for testing and evaluating unlearning systems. Lastly, this work identifies challenges and outlines future directions for improving the safety, reliability, and privacy compliance of machine unlearning, stressing the critical need for continued exploration in this area.